Circular No. 6172

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1995O IN ABELL 2218 J.-F. Le Borgne, Observatoire Midi-Pyrenées; B. Sanahuja, Univer-

sity of Barcelona; and S. Maddox, Royal Greenwich Observatory; report their discovery of an apparent supernova located at α = 16^{h}35^{m}56^{s}.121, δ = +66^{o}11′43′′.96 (equinox 2000.0), which is in the cluster Abell 2218 (cf. Abell et al. 1989, Ap.J. Suppl. 70, 1), and is 1′′ from arc 730 of Pello et al. (1992, A.Ap. 266, 6; note Fig. 7). The discovery was made on CCD images obtained on May 3.13 UT with the William Herschel Telescope (+ Low Dispersion Survey Spectrograph) at La Palma, with the magnitude estimated as R ∼ 22. SN 1995O does not appear on any of numerous frames obtained of this cluster since 1986, the last known images with no sign of the new star being J and K frames taken on 1994 Aug. 18 with the Canada-France-Hawaii Telescope (CFHT) at Mauna Kea. Low-dispersion spectrograms obtained by C. Soubiran (Observatoire de Bordeaux) and E. Friel (Maria Mitchell Observatory) on May 7.49 and 7.51 with the CFHT (+ multi-aperture spectrograph) show some signal between 450 and 900 nm, although the spectra are dominated by sky emission lines at wavelengths longer than 720 nm. Emissions at the redshift of Abell 2218 (z = 0.175) corresponding to features at ∼ 371.2, 464.5, 494.4, 551.1, and 584.8 nm (respectively) in the spectra of the type-Ia supernova 1989B (Wells et al. 1994, A.J. 108, 2233) were identified in the spectra of SN 1995O at 436.2, 545.8, 580.9, 647.5, and 687.1 nm. Likewise, the absorptions in SN 1989B at ∼ 486.2, 504.7, 569.2, and 602.0 nm were identified in SN 1995O at 571.3, 593.0, 668.8, and 707.3 nm.

COMET 71P/CLARK Visual magnitude estimates: Apr. 1.47 UT, 13.5 (A. Hale, Cloudcroft,

NM, 0.41-m reflector); 23.45, 12.9 (C. S. Morris, Pine Mountain Club, CA, 0.26-m reflector); 25.46, 12.4 (Hale); May 10.44, 11.6 (Hale).

NOVA AQUILAE 1995 Visual magnitude estimates: Apr. 21.43 UT, 10.5 (W. G. Dillon, Mis-

souri City, TX); May 1.17, 10.7 (A. Pereira, Cabo da Roca, Portugal); 6.115, 11.7 (P. Schmeer, Bischmisheim, Germany); 9.065, 11.7 (Schmeer).

1995 May 11 Daniel W. E. Green